# Conceptualization, Formal analysis, Investigation, Methodology, Software, Data curation, Validation, Visualization, Writing –original draft, Writing – review & editing

### v3

@@ 3.3 Common Feature Discriminator
Recognizing and leveraging these shared features offers a viable approach to establishing robust correspondences (Cheng and Cao, 2025) between LiDAR and camera detections of the same objects.
To achieve this, we propose the Common Feature Discriminator, a deep learning-based model designed to leverage deep learning’s superior feature extraction capabilities to learn and extract these shared features from LiDAR and camera data, thereby enabling effective object matching and correspondence identification. The first step of our approach is to detect and crop individual objects from each sensor’s output. To this end, we employ two established object detector tailored to LiDAR and camera data, respectively: 1) Camera-based object detection: We adopt YOLOv8 (Jocher, 2023) for robust 2D object detection in images. This produces bounding boxes around each detected object in the camera frame. 2) LiDAR-based object detection: We use an octree-based change detection algorithm (Strawlab, 2023) followed by a DBSCAN clustering step to segment point cloud regions corresponding to individual objects. This yields point clusters, each hypothesized to belong to a distinct object in the LiDAR frame. Because the LiDAR and camera frames are time-synchronized, each LiDAR cluster and camera bounding box within the same timestamp can be treated as candidate object detections from two complementary modalities.
Once objects have been cropped from the LiDAR and camera data, they are fed into our proposed Common Feature Discriminator. This module is responsible for determining whether an object in the camera image and an object in the LiDAR point cloud correspond to the same physical entity. To accomplish this, the Common Feature Discriminator learns and compares three key types of features—relative positions, appearance embeddings, and classification information—and fuses them to produce a final similarity decision, as outlined in Fig. 2:
  1. 1.
    LiDAR Feature Extraction: Each LiDAR object (i.e., a 3D point cluster) is passed to a LiDAR backbone (e.g., PointNet++), which encodes the local and global geometric structure of the object into a latent vector. A Classification Head outputs the object category (e.g., car, pedestrian)—accounting for the expectation that matching objects in LiDAR and camera views should share the same class—while an Embedding Head produces a 128-dimensional feature capturing the object’s 3D shape and reflection intensity. These two heads thus provide both semantic consistency checks (via classification) and geometric/reflective characteristics checks (via appearance embeddings). Let $\mathbf{X}_{L}$ denote the 3D point cluster associated with a LiDAR object. A PointNet++ backbone maps $\mathbf{X}_{L}$ to:
@@
    where $\mathbf{z}_{L}\in\mathbb{R}^{128}$ is the LiDAR embedding and $\hat{c}_{L}$ is the predicted class (e.g. car, truck, pedestrian).
  2. 2.
    Camera Feature Extraction: In parallel, each camera-cropped object (the pixels within its bounding box) is processed by an image backbone (e.g., ResNet). Similar to the LiDAR branch, this has a Classification Head that yields object categories, and an Embedding Head that provides a 128-dimensional representation of the geometric features (appearance embeddings) such as shape and structure. Let $\mathbf{I}_{C}$ be the pixel region representing a camera-detected object. A ResNet-based backbone maps $\mathbf{I}_{C}$ to:
@@
    providing the camera embedding $\mathbf{z}_{C}\in\mathbb{R}^{128}$ and the predicted semantic class $\hat{c}_{C}$.
  3. 3.
    Position Feature Extraction: Both LiDAR and camera objects also go through a Position Feature Extractor, which calculates the relative positions of objects within their respective sensor frames. This captures spatial alignment cues, allowing the model to verify if the apparent location of an object in the camera image aligns with the corresponding object’s location in the LiDAR point cloud. Each object’s 2D center, $(x,y)$ for LiDAR and $(u,v)$ for camera, yields:
    $\Delta\mathbf{p}\;=\;(u-x,\,v-y),$
    forming a relative position vector to capture spatial consistency across sensor views.
  4. 4.
    Feature Fusion and Matching: The three types of features—(i) Relative Positions derived by the Position Feature Extractor, (ii) Appearance Embeddings from the Embedding Heads, and (iii) Classification Information from the Classification Heads—are concatenated into a unified feature vector
    $\mathbf{f}\;=\;\big{[}\Delta\mathbf{p};\,\mathbf{z}_{L};\,\mathbf{z}_{C};\,\hat{c}_{L};\,\hat{c}_{C}\big{]},$
    which is then fed into a small feed-forward network (FFN) for binary classification (“Same” vs. “Differ”):
@@
    where $\sigma$ denotes the sigmoid activation. During training, a contrastive loss $\mathcal{L}_{\mathrm{ctr}}$ encourages embeddings of true-matching object pairs $(\hat{o}=1)$ to remain close while pushing non-matching pairs $(\hat{o}=0)$ apart, providing additional input for the classification decision.
@@
Algorithm 1 Iterative LiDAR–Camera Calibration Refinement
Input: $\text{Frames}\,F=\{(L_{i},C_{i})\}$ for $i=1\ldots T$trianglerightLiDAR and camera points for $T$ frames
[…]
     $N\in\mathbb{Z}$trianglerightNumber of frames to accumulate before recalibration
[…]
Output: $\text{CalibMatrix}^{*}$trianglerightRefined calibration matrix
[…]
3:for $i\in\{1,2,\dots,T\}$ do
4:     $\text{projPoints}\leftarrow\text{Project}(L_{i},\text{BestMatrix})$trianglerightProject LiDAR object centers onto the camera plane
[…]
8:     if $i\bmod N==0$ then
[…]
14:         end if
15:     end if
16:end for
17:$\text{CalibMatrix}^{*}\leftarrow\text{BestMatrix}$trianglerightFinal refined calibration matrix
18:return $\text{CalibMatrix}^{*}$
3.4 Homography-based Calibration Matrix Estimation
@@
To ensure robustness against erroneous matches, we employ the RANSAC algorithm (Cheng and Cao, 2023) to iteratively fit $\mathbf{H}$ while discarding outlier correspondences. Specifically, RANSAC randomly samples a small subset $\mathcal{C}_{s}\subset\mathcal{C}$ of correspondences to compute a candidate $\mathbf{H}_{s}$. It then evaluates $\mathbf{H}_{s}$ on the entire set $\mathcal{C}$ by measuring the reprojection error (e.g. $\mathcal{E}_{\text{AED}}$ or $\mathcal{E}_{\text{RMSE}}$), and repeats this process over multiple iterations. The matrix $\mathbf{H}$ yielding the largest inlier consensus (and thus the lowest average error) is ultimately selected.
@@
It is worth emphasizing that, unlike many existing LiDAR–camera calibration methods that attempt to utilize every LiDAR point, our approach relies solely on the centers of detected objects. We adopt this strategy for two main reasons (also as explained in Section 3.1). First, since the goal of calibration is to align LiDAR objects with camera objects, using object center points is already sufficient for establishing accurate correspondences; incorporating all LiDAR points does not provide any additional benefit for object association and can actually complicate the calibration matrix estimation process. Second, even though the calibration matrix is derived from object center points only, it can still be used to project the entire LiDAR point cloud onto the image plane. Moreover, this center-based approach naturally fits an object-level matching paradigm, especially considering that camera-detected objects lack corresponding point cloud data. By reducing the reliance on dense point sets and focusing on object centers, we gain more degrees of freedom to achieve a robust and flexible calibration outcome.
3.5 Iterative Refinement Process
@@
While relying on the Common Feature Discriminator to establish a coarse initial calibration matrix can provide a strong starting point, it may not perfectly match every corresponding object in LiDAR and camera data. In practice, leveraging additional data points (and thus increasing redundancy and field-of-view coverage) often improves both the accuracy and robustness of calibration estimation. To that end, we propose an iterative refinement procedure that successively updates the calibration matrix by incorporating newly discovered point correspondences across multiple frames. The process (as demonstrated in Algorithm 1) unfolds as follows:
  1. 1.
@@
    Initial Calibration: We begin by using the coarse calibration matrix obtained from coarse calibration as the initial matrix, i.e., $\mathbf{H}_{0}=\mathbf{H}_{\mathrm{coarse}}$. The LiDAR-camera point pairs identified by the Common Feature Discriminator during coarse calibration form an initial accumulated set $\mathcal{A}$.
  2. 2.
    LiDAR-to-Camera Projection: Let $\,\mathbf{H}_{\mathrm{best}}=\mathbf{H}_{0}$, for each incoming frame $\,(L_{i},\,C_{i})$, where $i\in\{1,\dots,T\}$, every LiDAR object center $\,(x_{j},\,y_{j})$ in $L_{i}$ is projected onto the camera plane using the current best matrix $\,\mathbf{H}_{\mathrm{best}}$:
@@
    We then compare each projected point $\bigl{(}\hat{u}_{j},\hat{v}_{j}\bigr{)}$ with the camera-detected objects in $C_{i}$.
  3. 3.
    Greedy Bipartite Matching: We then form a bipartite graph between the projected LiDAR points $\{\bigl{(}\hat{u}_{j},\hat{v}_{j}\bigr{)}\}$ and the camera-detected object centers $\{\bigl{(}u_{i},v_{i}\bigr{)}\}$ from $C_{i}$. To expand our pool of correspondences, a greedy bipartite matching (Besser and Poloczek, 2017) associates each LiDAR point with the closest camera object (if any) based on a distance measure $d\!\bigl{(}(\hat{u},\hat{v}),(u,v)\bigr{)}$.
  4. 4.
@@
    Correspondence Selection: Let $\mathcal{M}_{i}$ be the new candidate point pairs set from frame $i$. Only newly formed point pairs that fall within unoccupied or sufficiently distinct grid regions—determined using the proposed block-based sampling strategy (Fig. 5)—are retained. The resultant filtered set ~$\widetilde{\mathcal{M}}_{i}\subseteq\mathcal{M}_{i}$ is then added to the existing accumulated set $\,\mathcal{A}$:
    $\mathcal{A}\;\leftarrow\;\mathcal{A}\;\cup\;\widetilde{\mathcal{M}}_{i},$
    effectively expanding the coverage within the sensors’ field of view and enhancing the calibration robustness.
  5. 5.
@@
    Recalibration and Validation: After accumulating every $N=100$ frames (or another empirically chosen threshold) and gathering all associated point pairs, we re-estimate a new calibration matrix $\,\mathbf{H}_{\mathrm{new}}$ using the accumulated point pairs $\,\mathcal{A}$ and the homography calibration algorithm (Section 3.4). Formally,
@@
    $\mathbf{H}_{\mathrm{new}}=\mathrm{Recalibrate}\!\bigl{(}\mathcal{A}\bigr{)}\quad\text{via minimizing}\quad\sum_{\bigl{(}x,y\bigr{)}\leftrightarrow\bigl{(}u,v\bigr{)}\,\in\,\mathcal{A}}\varphi\!\Bigl{(}\mathbf{H},\,(x,y),\,(u,v)\Bigr{)},$
    where $\,\varphi(\cdot)$ is the chosen reprojection error function (e.g. $\,\mathcal{E}_{\mathrm{AED}}$, $\,\mathcal{E}_{\mathrm{RMSE}}$). If the new matrix provides no improvement in terms of reprojection error (i.e., projected LiDAR points are not better aligned with camera objects), it is discarded; otherwise, it updates the best matrix $\,\mathbf{H}_{\mathrm{best}}\leftarrow\mathbf{H}_{\mathrm{new}}$.
  6. 6.
    Iterate: We repeat Steps 2–5 for all subsequent frames $i+1,i+2,\dots$ until reaching the final time step $T$. This iterative loop gradually refines the calibration matrix by incorporating newly validated point correspondences.
By systematically incorporating additional correspondences at each iteration, this optimization loop converges toward a more robust calibration matrix. It maintains the practical advantages of the initial deep learning–based matching while progressively enhancing accuracy through redundancy and extended spatial coverage. Moreover, its iterative nature naturally accommodates runtime changes in the environment, thus helping to mitigate potential decalibration over long-term operation.
@@
It is worth noting that we opt for a greedy bipartite matching (Besser and Poloczek, 2017) approach rather than the more popular Hungarian algorithm for several practical reasons. Ideally, each LiDAR detection would correspond to exactly one camera detection, and bipartite graph matching would produce a one-to-one mapping that minimizes the overall matching cost. However, real-world conditions deviate from this ideal scenario: variations in field of view and detection capabilities can lead to certain objects being detected by only one sensor. For example, LiDAR may capture distant objects outside the camera’s range, whereas a camera may pick up small or reflective objects that the LiDAR cannot reliably detect. Given these discrepancies, the goal of bipartite graph matching is to identify the best subset of matching pairs, without forcing all detections from both sensors to be paired. Greedy bipartite matching is well-suited to this task, as it prioritizes finding and accumulating the lowest-cost matches while allowing some objects to remain unmatched if no suitable pair exists. In contrast, the Hungarian algorithm aims for an optimal, one-to-one, and complete assignment—i.e., pairing every detection from both sensors—an assumption that does not hold in many real-world LiDAR–camera detections. Such forced one-to-one pairings can degrade matching quality when unmatchable objects are forced to pair with unrelated detections.
Table 1: Performance of Common Feature Discriminator
@@
3.6 Attention-based Refinement Process
While homography-based calibration can provide a reasonable initial solution, it relies on planar assumptions and may introduce non-negligible errors in real-world environments with complex depth variations. To address these limitations and further refine the calibration, we adopt an attention-based deep learning model that produces a correction matrix $\mathbf{H}^{\Delta}$, as illustrated in Fig. 4. Once trained, the refined calibration is computed as

### v4

@@ III-C Common Feature Discriminator
Recognizing and leveraging these shared features offers a viable approach to establishing robust correspondences [45] between LiDAR and camera detections of the same objects.
To achieve this, we propose the Common Feature Discriminator, a deep learning–based model that leverages advanced feature extraction to learn and extract shared features from LiDAR and camera data, thereby enabling effective object matching and correspondence identification. The first step is to detect and crop individual objects from each sensor’s output. For camera-based object detection, we adopt YOLOv8 [46] to robustly detect objects in images and generate bounding boxes around them. In parallel, LiDAR-based object detection is performed using an octree-based change detection algorithm [47] followed by DBSCAN clustering, which segments the point cloud into clusters, each hypothesized to belong to a distinct object. Since the LiDAR and camera frames are time-synchronized, each LiDAR cluster and corresponding camera bounding box at the same timestamp can be treated as candidate detections from complementary modalities.
Once the objects are cropped from both sensor outputs, they are fed into the Common Feature Discriminator, whose task is to determine whether an object in a camera image and an object in a LiDAR point cloud correspond to the same physical entity. To this end, the discriminator learns and compares three key types of features: relative positions, appearance embeddings, and classification information. In the LiDAR branch, each 3D point cluster $\mathbf{X}_{L}$ is processed by a LiDAR backbone (e.g., PointNet++) that encodes its local and global geometric structure into a latent vector. A classification head then outputs the object category while an embedding head produces a 128-dimensional feature, yielding
@@
where $\mathbf{z}_{L}\in\mathbb{R}^{128}$ represents the LiDAR embedding and $\hat{c}_{L}$ the predicted class.
Simultaneously, each camera-cropped object (the pixels within its bounding box) is processed by an image backbone (e.g., ResNet), which outputs both a 128-dimensional appearance embedding and a classification result:
@@
where $\mathbf{z}_{C}\in\mathbb{R}^{128}$ is the camera embedding and $\hat{c}_{C}$ the predicted semantic class. In addition, both LiDAR and camera objects are passed through a position feature extractor that computes the relative positions from their respective 2D centers, $(x,y)$ for LiDAR and $(u,v)$ for the camera, resulting in a relative position vector
$\Delta\mathbf{p}=(u-x,\,v-y).$
These three types of features—relative positions, appearance embeddings, and classification information—are then concatenated into a unified feature vector:
$\mathbf{f}\;=\;\bigl{[}\Delta\mathbf{p};\,\mathbf{z}_{L};\,\mathbf{z}_{C};\,\hat{c}_{L};\,\hat{c}_{C}\bigr{]},$
which is fed into a small feed-forward network (FFN) for binary classification (“Same” vs. “Differ”):
@@
where $\sigma$ denotes the sigmoid activation. During training, a contrastive loss $\mathcal{L}_{\mathrm{ctr}}$ encourages the embeddings of true matching pairs ($\hat{o}=1$) to be close, while pushing non-matching pairs ($\hat{o}=0$) apart.
@@
III-D Homography-based Calibration Matrix Estimation
@@
To ensure robustness against erroneous matches, we employ the RANSAC algorithm [18] to iteratively fit $\mathbf{H}$ while discarding outlier correspondences. Specifically, RANSAC randomly samples a small subset $\mathcal{C}_{s}\subset\mathcal{C}$ of correspondences to compute a candidate $\mathbf{H}_{s}$. It then evaluates $\mathbf{H}_{s}$ on the entire set $\mathcal{C}$ by measuring the reprojection error (e.g. $\mathcal{E}_{\text{AED}}$ or $\mathcal{E}_{\text{RMSE}}$), and repeats this process over multiple iterations. The matrix $\mathbf{H}$ yielding the largest inlier consensus (and thus the lowest average error) is ultimately selected.
@@
It is worth emphasizing that, unlike many existing LiDAR–camera calibration methods that attempt to utilize every LiDAR point, our approach relies solely on the centers of detected objects. We adopt this strategy for two main reasons (also as explained in Section III-A). First, since the goal of calibration is to align LiDAR objects with camera objects, using object center points is already sufficient for establishing accurate correspondences; incorporating all LiDAR points does not provide any additional benefit for object association and can actually complicate the calibration matrix estimation process. Second, even though the calibration matrix is derived from object center points only, it can still be used to project the entire LiDAR point cloud onto the image plane. Moreover, this center-based approach naturally fits an object-level matching paradigm, especially considering that camera-detected objects lack corresponding point cloud data. By reducing the reliance on dense point sets and focusing on object centers, we gain more degrees of freedom to achieve a robust and flexible calibration outcome.
Algorithm 1 Iterative LiDAR–Camera Calibration Refinement
Input: $F=\{(L_{i},C_{i})\}$ for $i=1\ldots T$trianglerightLiDAR and camera points for $T$ frames
     $H_{0}$trianglerightInitial calib. matrix from CFD
     $n\in\mathbb{Z}$trianglerightFrames to accumulate before recalib.
     $Bsz$trianglerightBlock size for sampling
Output: $H^{*}$trianglerightRefined calibration matrix
1:$H_{best}\leftarrow H_{0}$trianglerightSet initial matrix as best
2:$A\leftarrow[\ ]$trianglerightList to store pairs
3:for $i\in\{1,2,\dots,T\}$ do
4:     $P^{L}_{i}\leftarrow proj(pts(L_{i}),H_{best})$trianglerightProject LiDAR points onto the camera plane
5:     $m\leftarrow GBMatch(P^{L}_{i},pts(C_{i}))$trianglerightGreedy Match points
6:     $s\leftarrow BBSample(m,Bsz)$trianglerightBlock-based sampling
7:     $A.\text{append}(s)$trianglerightAccumulate new pairs
8:     if $i\bmod n==0$ then
9:         $H_{new}\leftarrow Recalib(A)$trianglerightRecalibrate using accumulated pairs
10:         $e_{old}\leftarrow RepErr(H_{best},A)$
11:         $e_{new}\leftarrow RepErr(H_{new},A)$
12:         if $e_{new}<e_{old}$ then
13:              $H_{best}\leftarrow H_{new}$trianglerightUpdate matrix if error reduces
14:         end if
15:     end if
16:end for
17:$H^{*}\leftarrow H_{best}$trianglerightFinal refined calibration matrix
18:return $H^{*}$
III-E Iterative Refinement Process
@@
While relying on the Common Feature Discriminator to establish a coarse initial calibration matrix provides a strong starting point, it may not perfectly match every corresponding object between LiDAR and camera data. In practice, leveraging additional data points—thereby increasing redundancy and expanding field-of-view coverage—often improves both the accuracy and robustness of the calibration. To this end, we propose an iterative refinement procedure (as demonstrated in Algorithm 1) that successively updates the calibration matrix by incorporating newly discovered point correspondences across multiple frames.
@@
We begin with the coarse calibration matrix, denoted as $\mathbf{H}_{0}=\mathbf{H}_{\mathrm{coarse}}$, and use the LiDAR–camera point pairs identified during the coarse calibration (by using the Common Feature Discriminator) to form an initial accumulated set $\mathcal{A}$. For each incoming frame $(L_{i},C_{i})$, where $i\in\{1,\dots,T\}$, every LiDAR object center $(x_{j},y_{j})$ in $L_{i}$ is projected onto the camera plane using the current best calibration matrix $\mathbf{H}_{\mathrm{best}}$ as follows:
@@
Each projected point $(\hat{u}_{j},\hat{v}_{j})$ is then compared with the camera-detected object centers in $C_{i}$. A greedy bipartite matching algorithm [48] is used to associate each projected LiDAR point with its nearest camera detection (if one exists) based on a distance measure $d\bigl{(}(\hat{u},\hat{v}),(u,v)\bigr{)}$. Let $\mathcal{M}_{i}$ be the associated candidate point pairs set from frame $i$. Only candidate point pairs that fall within unoccupied or sufficiently distinct grid regions—determined by our block-based sampling strategy (Fig. 4)—are retained to form a filtered set ~$\widetilde{\mathcal{M}}_{i}\subseteq\mathcal{M}_{i}$, which is then incorporated into the accumulated set $\mathcal{A}$ via
$\mathcal{A}\leftarrow\mathcal{A}\cup\widetilde{\mathcal{M}}_{i}.$
@@
After accumulating data from every $N=100$ frames (or another empirically chosen threshold), a new calibration matrix $\mathbf{H}_{\mathrm{new}}$ is re-estimated from the set $\mathcal{A}$ using the homography calibration algorithm described in Section III-D:
@@
$\mathbf{H}_{\mathrm{new}}$ $=\mathrm{Recalibrate}\bigl{(}\mathcal{A}\bigr{)}\quad\text{via minimizing}$
$\quad\sum_{(x,y)\leftrightarrow(u,v)\in\mathcal{A}}\varphi\bigl{(}\mathbf{H},(x,y),(u,v)\bigr{)},$
where $\varphi(\cdot)$ denotes the chosen reprojection error function (e.g., $\mathcal{E}_{\mathrm{AED}}$ or $\mathcal{E}_{\mathrm{RMSE}}$). If $\mathbf{H}_{\mathrm{new}}$ results in a reduced reprojection error, it replaces the current best matrix, i.e., $\mathbf{H}_{\mathrm{best}}\leftarrow\mathbf{H}_{\mathrm{new}}$. This process is iterated for each subsequent frame until reaching the final time step $T$, gradually refining the calibration matrix by incorporating newly validated point correspondences.
@@
By systematically incorporating additional correspondences at each iteration, this optimization loop converges toward a more robust calibration matrix. It maintains the practical advantages of the initial deep learning–based matching while progressively enhancing accuracy through redundancy and extended spatial coverage. Moreover, its iterative nature naturally accommodates runtime changes in the environment, thus helping to mitigate potential decalibration over long-term operation. Notably, we opt for a greedy bipartite matching [48] approach rather than the more popular Hungarian algorithm for several practical reasons. Ideally, each LiDAR detection would correspond to exactly one camera detection, and bipartite graph matching would produce a one-to-one mapping that minimizes the overall matching cost. However, real-world conditions deviate from this ideal scenario: variations in field of view and detection capabilities can lead to certain objects being detected by only one sensor. For example, LiDAR may capture distant objects outside the camera’s range, whereas a camera may pick up small or reflective objects that the LiDAR cannot reliably detect. Given these discrepancies, the goal of bipartite graph matching is to identify the best subset of matching pairs, without forcing all detections from both sensors to be paired. Greedy bipartite matching is well-suited to this task, as it prioritizes finding and accumulating the lowest-cost matches while allowing some objects to remain unmatched if no suitable pair exists. In contrast, the Hungarian algorithm aims for an optimal, one-to-one, and complete assignment—i.e., pairing every detection from both sensors—an assumption that does not hold in many real-world LiDAR–camera detections. Such forced one-to-one pairings can degrade matching quality when unmatchable objects are forced to pair with unrelated detections.
TABLE I: Performance of Common Feature Discriminator
@@
III-F Attention-based Refinement Process
While homography-based calibration yields a reasonable initial solution, its reliance on planar assumptions often introduces significant errors in real-world environments characterized by complex depth variations. To overcome these limitations and further refine the calibration, we introduce an attention-based deep learning model that produces a correction matrix $\mathbf{H}^{\Delta}$ (see Fig. 3). The refined calibration is computed as